# HAFO: A Force-Adaptive Control Framework for Humanoid Robots in Intense Interaction Environments

### v3

@@ Abstract
Reinforcement learning (RL) controllers have made impressive progress in humanoid locomotion and light-weight object manipulation. However, achieving robust and precise motion control with intense force interaction remains a significant challenge. To address these limitations, this paper proposes HAFO, a dual-agent reinforcement learning framework that concurrently optimizes both a robust locomotion strategy and a precise upper-body manipulation strategy via coupled training in environments with external disturbances. The external pulling disturbances are explicitly modeled using a spring-damper system, allowing for fine-grained force control through manipulation of the virtual spring. In this process, the reinforcement learning policy autonomously generates a disturbance-rejection response by utilizing environmental feedback. Furthermore, HAFO employs an asymmetric Actor-Critic framework in which the Critic network’s access to privileged external forces guides the actor network to acquire generalizable force adaptation for resisting external disturbances. The experimental results demonstrate that HAFO achieves whole-body control for humanoid robots across diverse force-interaction environments, delivering outstanding performance in load-bearing tasks and maintaining stable operation even under rope suspension state.
@@
Force-adaptive Control , Whole body control , Reinforcement learning
@@
[4]organization=Shanghai AI Laboratory,city=Shanghai, postcode=200030, country=China
@@ 1 Introduction
In recent years, learning-based control solutions have made significant progress in the field of whole-body control for humanoid robots [1, 2, 3, 4, 5, 6]. By leveraging large-scale parallel simulations and domain randomization, policy networks can effectively learn to handle environmental uncertainties, thus enabling zero-shot transfer to real robots. However, RL controllers often overlook or oversimplify interactive forces. To address the practical needs of humanoid robots in tasks such as heavy-load manipulation and high-altitude operations, precise and robust motion control is crucial in environments with strong external forces and dynamic changes.
@@
When humanoid robots encounter strong collisions or physical contact, reinforcement learning (RL) controllers often exhibit instability. Specifically, current RL methods primarily rely on high-level commands, such as movement velocity, root height, or predefined trajectories, which may mitigate some random disturbances. However, when exposed to human intervention or significant environmental contact, the absence of explicit modeling of contact dynamics often hinders the robot’s ability to maintain balance, potentially leading to safety hazards [7, 8]. For example, in high-altitude operations, ropes are required for support, but their interference can significantly undermine the robot’s stability. In heavy-load scenarios, large payloads significantly impair balance, placing considerable demands on the load-bearing capacity.
@@
To enhance the stability of control strategies under human intervention or significant external disturbances, this paper introduces HAFO, a force-adaptive control framework based on explicit dynamic modeling. HAFO employs a dual-agent strategy that decouples the control of the upper and lower bodies. The lower-body strategy focuses on robust locomotion to maintain gait stability under strong external disturbances. The upper-body strategy emphasizes precise and real-time upper-body manipulation under external force disturbances. We utilize a spring-damper mechanical system to model external pulling disturbances. Meanwhile, we progressively increase external forces through a curriculum-learning schedule and randomize the point of force application, systematically enhancing the strategy’s generalization to various pulling disturbances and different hand loads. The contributions of this paper are as follows:
  1. 1)
@@
    A dual-agent reinforcement learning framework is proposed to facilitate the collaborative evolution of the upper and lower bodies in the presence of large external disturbances. The lower-body strategy focuses on ensuring robust locomotion, and the upper-body strategy focuses on precise manipulation. Both strategies share a common state space. Through a coupling-adversarial mechanism, the stability of the entire body system under disturbances is significantly improved, enhancing the robot’s control robustness against multi-source external disturbances.
  2. 2)
@@
    A spring-damping dynamic model is developed to explicitly model the external pulling force as equivalent spring-damper forces acting on the robot’s body. During model training, a carefully designed progressive curriculum learning strategy is employed, gradually increasing the spring-damper forces to enable the policy to adapt progressively to external disturbances. The reinforcement learning (RL) policy autonomously evolves force-adaptive control modes for ground locomotion and aerial suspension, spontaneously generating mode-switching behaviors based on environmental feedback, without relying on explicit state machines or predefined switching logic.
  3. 3)
@@
    Experimental results demonstrate that the HAFO policy enables adaptive force control across various environments, exhibiting stable and precise motion capabilities under single or dual-hand loads. Moreover, HAFO is the first locomotion-control strategy to achieve stable operation and safe startup under rope-suspension conditions, marking preliminary exploratory work for humanoid robot applications in high-altitude environments.
@@ 2.1 Whole-body control based on reinforcement learning
Conventional dynamic modeling methods have shown impressive control performance in structured scenarios for legged robots [9, 10, 11, 12, 13, 14, 15, 16, 17]. However, due to the intrinsic under-actuation of humanoid robots and the variability of real-world environments, it is extremely difficult to establish an accurate dynamics model. Recently, several RL-based solutions have provided new insights for whole-body control of humanoid robots [1, 2, 3, 4, 6, 7, 8, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32]. Common paradigms include Lower-RL-Upper-IK and Integrated Whole-body RL, which address the challenge of whole-body motion control from different perspectives.
@@
In terms of Lower-RL-Upper-IK paradigm, Ben et al. [18] proposed the Homie, which employs curriculum learning to dynamically adjust the motion tracking threshold, allowing the policy to progressively adapt to high-difficulty motions during training. Additionally, the Homie controller develops an isomorphic exoskeleton cockpit to achieve precise whole-body control of the robot. By combining large-scale human motion data with reinforcement learning, Cheng et al. [19] encourage the robot’s upper body to imitate the reference motion, while relaxing the imitation constraints on the legs, allowing the robot to achieve expressive movements. The Mobile-TeleVision proposed by Lu et al. [23] introduces the predictive motion prior (PMP), which encodes multi-modal motion features through conditional variational autoencoder (CVAE) to guide the reinforcement learning strategy generates dynamic motions. For the Integrated Whole-body RL paradigm, Xie et al. [28] proposed a KungfuBot, which incorporates foot-ground contact states and a motion-phase signal to accurately imitate highly dynamic wushu movements. He et al. [2] introduced the OmniH2O teleoperation system, which integrates a teacher-student distillation framework to transfer knowledge from a privileged teacher policy to the deployed policy, enabling efficient and stable motion control of the robot in real-world environments. He et al. [21] proposed ASAP, which significantly enhances a robot’s adaptability in complex environments by leveraging a pretrained motion tracking policy with a residual action model trained using real-world data. Liao et al. [27] designed a scalable and high-quality motion tracking framework called BeyondMimic, which transforms kinematic references into robust, highly dynamic robot motions. Through distillation, it integrates the trained motions into a single policy, enabling the flexible combination of multiple skills during testing and achieving goal-driven control using simple cost functions.
@@
Although whole-body motion control for humanoid robots has made remarkable progress in the research mentioned above, the existing methods still face many challenges. On the one hand, the Lower-RL-Upper-IK paradigm employs an open-loop control mode for the upper body, which struggles to perceive and adapt to external interference in real time. On the other hand, the Integrated paradigm controls whole-body joints through a single controller, but due to the weak task goal correlation between upper-body manipulation and lower-body walking, it is prone to overfitting.
@@ 2.2 Force-adaptive control
Portela [33] et al. first introduced the force-adaptive control objective for quadruped robots, which significantly improved the force interaction capability of the end effector. However, the model requires explicit switching between position mode and force mode, and assumes the system is in a quasi-static state, making it difficult to handle dynamic and unknown disturbances. For humanoid robots, their inherent characteristics, such as a high center of gravity and a small support base, make force-adaptive control particularly complex [34, 35, 36]. Li et al. [37] introduced a force-adaptive torso-tilt reward, which maintains balance and effectively transmits force through torso tilting and center-of-gravity adjustment, generating greater pulling forces in high-intensity force interaction tasks. Zhang et al. [6] designed a 3D force curriculum, which maximizes the force adaptability while ensuring the feasibility of joint torque, achieving high-intensity force interactions. Kuo et al. [38] proposed a hybrid control method combining the Artificial Rabbit Optimization (ARO) algorithm with an MLP neural network. Through multi-angle impact experiments on flat and sloped terrain, they verified that the method can quickly adapt to external forces and effectively maintain balance. Existing model-based methods rely on precise modeling of pre-planned force trajectories, which restricts their generalization capability in unknown environments, rendering them inadequate for adapting to dynamically changing conditions [39]. In contrast, estimation-based approaches [40] are only applicable under quasi-static conditions and fail to effectively address force adaptive control in dynamic loco-manipulation tasks.
Although existing studies have explored force adaptive control for humanoid robots, the range of scenarios explored so far remains relatively limited. In this paper, we propose a force-adaptive whole-body control strategy (HAFO), which endows humanoid robots with force-adaptive capabilities in challenging conditions, such as heavy loads and pulling disturbances.
@@ 3 Method
Humanoid robot motion control requires coordinated movement of whole-body joints. Traditional whole-body control methods rely on a single strategy to output target joint angles for the entire body. However, the lack of strong coupling between upper and lower bodies tasks leads to inefficient sample utilization, making the policy prone to overfitting. To maximize the exploration efficiency of reinforcement learning, this paper proposes the HAFO architecture shown in Figure. 1, which employs a dual-agent reinforcement learning framework to drive the co-evolution of upper and lower bodies control strategies. Additionally, during training, various dynamic disturbances are explicitly introduced to ensure stable force-adaptive control.
@@ 3.1 Dual-agent learning framework
We decompose the total degrees of freedom (DOFs) of the humanoid into upper and lower bodies, such that $n_{\mathrm{dof}}=\mathit{\text{Upper}}_{\mathrm{\text{dof}}}+\mathit{\text{Lower}}_{\mathrm{dof}}$, where $\mathit{\text{Upper}}_{\text{dof}}$ is controlled by upper-body strategy $\pi_{\text{Upper}}\left(\mathrm{a}_{\text{Upper }}\mid\mathrm{s}\right)$, $\mathit{\text{Lower}}_{\text{dof}}$ is controlled by the lower-body strategy $\pi_{\text{lower }}\left(\mathrm{a}_{\text{lower }}\mid\mathrm{s}\right)$. The dual-agent reinforcement learning strategy is trained using the Proximal Policy Optimization (PPO) algorithm [41] and shares the same state space ${s}_{{t}}=\left[{s}_{{t}}^{\text{prop }},\mathrm{c}^{l},\mathrm{~g}^{{u}}\right]$, allowing both agents to consider information from the entire body and achieve coordinated control. Among them, the ${s}_{\mathrm{t}}^{\text{prop }}=\left[{q}_{{t}},\dot{q}_{t},\omega_{t},gv_{t},a_{t-1}\right]$ represents the body perception information of the robot, the $c^{l}=\left[\hat{v}_{x,t},\hat{v}_{y,t},\hat{\omega}_{yaw},t\right]$ represents the command information for the lower-body strategy, and the ${g}^{{u}}\in\mathbb{R}^{14}$ represents the command information for the upper-body strategy.
The dual-agent policy is formulated as a dual-objective optimization problem. The value functions $\left(V^{l},V^{u}\right)$ corresponding to the upper and lower body policies $\left(\pi^{l},\pi^{u}\right)$, and the rewards obtained can be represented as $\left(r^{l}\left(s,a^{l},a^{u}\right),r^{u}\left(\mathrm{~s},\mathrm{a}^{1},\mathrm{a}^{\mathrm{u}}\right)\right)$. The dual policies are optimized by maximizing the cumulative rewards [42],
@@
3.2 Robust lower-body movement under external forces
Robust locomotion of humanoid robots remains a significant challenge, primarily due to the complexity of the floating base and the instability caused by high inertia. This difficulty is further compounded when payloads are applied to the upper body, increasing the overall inertia and introducing additional modeling uncertainties, which complicate the control design.
For the lower-body strategy $\pi^{l}:C^{l}\times S\mapsto A^{l}$, various external disturbances are introduced during the training process to enhance the robot’s robustness, including adversarial upper-body perturbations and high-frequency random pushing disturbances.
3.3 Upper-body motion generation based on imitation learning
The demonstration data for the upper-body policy are derived from the AMASS human-motion dataset [43] after feasibility filtering. Following PHC [24], human poses are retargeted to humanoid poses, and filtering is applied to suppress jitter and discontinuities. Instead of using open-loop replay, we train a reinforcement learning policy for the upper body that adapts to changing environments, avoids self-collisions, and improves whole-body coordination. The upper-body policy $\pi^{\mathrm{u}}:C^{{u}}\times S\mapsto A^{{u}}$ is trained in conjunction with the lower-body policy $\pi^{l}:C^{l}\times S\mapsto A^{l}$, and the state space of the upper-body strategy includes the state of the entire body. To enhance the load capacity of the humanoid robots, disturbances are applied to the robot’s end-effectors in the form of random 3D spatial forces [6].

### v4

@@ Abstract
Reinforcement learning (RL) controllers have made impressive progress in humanoid locomotion and light-weight object manipulation. However, achieving robust and precise motion control with intense force interaction remains a significant challenge. To address these limitations, this paper proposes HAFO, a dual-agent reinforcement learning framework that concurrently optimizes both a robust locomotion strategy and a precise upper-body manipulation strategy via coupled training. We employ a constrained residual action space to improve dual-agent training stability and sample efficiency. The external tension disturbances are explicitly modeled using a spring-damper system, allowing for fine-grained force control through manipulation of the virtual spring. In this process, the reinforcement learning policy autonomously generates a disturbance-rejection response by utilizing environmental feedback. The experimental results demonstrate that HAFO achieves whole-body control for humanoid robots across diverse force-interaction environments using a single dual-agent policy, delivering outstanding performance under load-bearing and thrust-disturbance conditions, while maintaining stable operation even under rope suspension state.
@@
Humanoid robot , Force-adaptive control , Whole-body control , Spring-damper system , Dual-agent reinforcement learning
@@
[4]organization=Shanghai Artificial Intelligence Laboratory,city=Shanghai, postcode=200030, country=China
@@ 1 Introduction
In recent years, learning-based control solutions have made significant progress in the field of whole-body control for humanoid robots [11, 13, 31, 43, 9, 48]. By leveraging large-scale parallel simulations and domain randomization, policy networks can effectively learn to handle environmental uncertainties, thus enabling zero-shot transfer to real robots. However, current RL controllers often overlook or oversimplify interactive forces. To address the practical needs of humanoid robots in tasks such as heavy-load manipulation and high-altitude operations, achieving precise and robust motion control is crucial in environments with strong external force disturbances.
@@
When humanoid robots encounter strong collisions or physical contact, reinforcement learning controllers often exhibit instability. Specifically, current RL methods primarily rely on high-level commands, such as movement velocity, root height, or predefined trajectories, which may mitigate some random disturbances. However, when exposed to human intervention or significant environmental contact, the absence of explicit modeling of contact dynamics often hinders the robot’s ability to maintain balance, potentially leading to safety hazards [35, 42]. For example, in high-altitude operations, ropes are required for support, but their interference can significantly undermine the robot’s stability. In heavy-load scenarios, large payloads significantly impair balance, placing considerable demands on the load-bearing capacity.
@@
To enhance the stability of control strategies under human intervention or significant external disturbances, this paper introduces HAFO, a force-adaptive control framework based on explicit dynamic modeling. HAFO employs a dual-agent strategy that decouples the control of the upper and lower bodies. The lower-body strategy focuses on robust locomotion to maintain gait stability under strong external disturbances. The upper-body strategy emphasizes precise and real-time upper-body manipulation under external force disturbances. We utilize a spring-damper mechanical system to model external tension disturbances. Meanwhile, we progressively increase external forces through a curriculum-learning schedule and randomize the point of force application, systematically enhancing the strategy’s generalization to various force disturbances. The contributions of this paper are as follows:
  1. (1)
@@
    A dual-agent reinforcement learning framework is proposed to facilitate the collaborative evolution of the upper and lower bodies in the presence of large external disturbances. The lower-body strategy focuses on ensuring robust locomotion, and the upper-body strategy focuses on precise manipulation. For upper-body training, we introduce a constrained residual action space, which significantly enhances the stability of the two agents during training. Through adversarial training, we achieve stable and efficient coordinated whole body control on the Unitree G1, and successfully extend it to the full-sized humanoid robot Unitree H1-2, further demonstrating its robustness and strong scalability.
  2. (2)
@@
    A spring-damping dynamic model is developed to explicitly model the external tension force as equivalent spring-damper forces acting on the robot’s body. During model training, a progressive curriculum learning strategy is employed, gradually increasing the spring-damper forces to enable the policy to adapt progressively to external disturbances. The reinforcement learning policy autonomously evolves force-adaptive control modes for ground locomotion and aerial suspension, spontaneously generating mode-switching behaviors based on environmental feedback, without relying on explicit state machines or predefined switching logic.
  3. (3)
@@
    Experimental results demonstrate that HAFO achieves force-adaptive control for humanoid robots under diverse conditions such as heavy load, thrust disturbance and rope suspension using a single dual-agent policy, and exhibits stable and precise motion capabilities under single or dual-hand loads. Moreover, HAFO is the first locomotion-control strategy to achieve stable operation and safe startup under rope-suspension conditions, marking preliminary exploratory work for humanoid robot applications in high-altitude environments.
@@ 2.1 Whole-body control based on reinforcement learning
Conventional dynamic modeling methods have shown impressive control performance in structured scenarios for legged robots [3, 36, 39, 8, 16, 6, 44, 7, 40]. However, due to the intrinsic under-actuation of humanoid robots and the variability of real-world environments, it is extremely difficult to establish an accurate dynamics model. Recently, several RL-based solutions have provided new insights for whole-body control of humanoid robots [11, 13, 31, 43, 48, 35, 42, 2, 5, 10, 12, 21, 23, 24, 34, 37, 22, 41, 45, 46, 47, 17]. Common paradigms include lower-RL-upper-IK and integrated-whole-body-RL, which address the challenge of whole-body motion control from different perspectives.
@@
In terms of lower-RL-upper-IK paradigm, Ben et al. [2] proposed the Homie, which employs curriculum learning to dynamically adjust the motion tracking threshold, allowing the policy to progressively adapt to high-difficulty motions during training. By combining large-scale human motion data with reinforcement learning, Cheng et al. [5] encourage the robot’s upper body to imitate the reference motion, while relaxing the imitation constraints on the legs, allowing the robot to achieve expressive movements. The Mobile-TeleVision proposed by Lu et al. [23] introduces the predictive motion prior (PMP), which encodes multi-modal motion features through conditional variational autoencoder (CVAE) to guide the reinforcement learning strategy generates dynamic motions. For the integrated-whole-body-RL paradigm, Xie et al. [41] proposed a KungfuBot, which incorporates foot-ground contact states and a motion-phase signal to accurately imitate highly dynamic wushu movements. He et al. [13] introduced the OmniH2O teleoperation system, which integrates a teacher-student distillation framework to transfer knowledge from a privileged teacher policy to the deployed policy, enabling efficient and stable motion control of the robot in real-world environments. He et al. [12] proposed ASAP, which significantly enhances a robot’s adaptability in complex environments by leveraging a pretrained motion tracking policy with a residual action model trained using real-world data. Liao et al. [22] designed a scalable and high-quality motion tracking framework called BeyondMimic, which transforms kinematic references into robust, highly dynamic robot motions. Through distillation, it integrates the trained motions into a single policy, enabling the flexible combination of multiple skills during testing and achieving goal-driven control using simple cost functions.
@@
Although whole-body motion control for humanoid robots has made remarkable progress in the research mentioned above, the existing methods still face many challenges. On the one hand, the lower-RL-upper-IK paradigm employs an open-loop control mode for the upper body, which struggles to perceive and adapt to external interference in real time. On the other hand, the integrated-whole-body-RL paradigm controls whole-body joints through a single controller, but due to the weak task goal correlation between upper-body manipulation and lower-body walking, it is prone to overfitting.
@@ 2.2 Force-adaptive control
Portela [30] et al. first introduced the force-adaptive control objective for quadruped robots, which significantly improved the force interaction capability of the end effector. However, the model requires explicit switching between position mode and force mode, and assumes the system is in a quasi-static state, making it difficult to handle dynamic and unknown disturbances. For humanoid robots, their inherent characteristics, such as a high center of gravity and a small support base, make force-adaptive control particularly complex [27, 1, 4]. Li et al. [19] introduced a force-adaptive torso-tilt reward, which maintains balance and effectively transmits force through torso tilting and center-of-gravity adjustment, generating greater pulling forces in high-intensity force interaction tasks. Zhang et al. [48] designed a force curriculum, which maximizes the force adaptability while ensuring the feasibility of joint torque, achieving high-intensity force interactions. Kuo et al. [18] proposed a hybrid control method combining the Artificial Rabbit Optimization (ARO) algorithm with an MLP neural network. Through multi-angle impact experiments on flat and sloped terrain, the method can quickly adapt to external forces and effectively maintain balance. Existing model-based methods rely on precise modeling of pre-planned force trajectories, which restricts their generalization capability in unknown environments, rendering them inadequate for adapting to dynamically changing conditions [20]. In contrast, estimation-based approaches [28] are only applicable under quasi-static conditions and fail to effectively address force adaptive control in dynamic loco-manipulation tasks.
Although some works have studied force adaptive control for humanoid robots, the range of scenarios explored so far remains relatively limited. In this paper, we propose a force-adaptive whole-body control strategy (HAFO), which endows humanoid robots with force-adaptive capabilities in challenging conditions, such as heavy loads, thrust disturbance, and tension disturbances using a single dual-agent policy.
@@ 3 Method
Humanoid robot motion control requires coordinated movement of whole-body joints. Traditional whole-body control methods rely on a single strategy to output target joint angles for the entire body. However, the lack of strong coupling between upper and lower bodies tasks leads to inefficient sample utilization, making the policy prone to overfitting. To maximize the exploration efficiency of reinforcement learning, this paper proposes HAFO architecture shown in Fig. 1, which employs a dual-agent reinforcement learning framework to drive the co-evolution of upper and lower bodies control strategies. During training, various dynamic disturbances are explicitly introduced to ensure stable control.
@@ 3.1 Dual-agent learning framework
We decompose the total degrees of freedom (DOFs) of the humanoid into upper and lower bodies, such that $n_{\text{dof }}=n_{\text{dof }}^{\text{upper }}+n_{\text{dof }}^{\text{lower }}$, where $n_{\text{dof }}^{\text{upper }}$ is controlled by upper-body strategy $\pi_{\text{upper}}\left(a_{\text{upper }}\mid\mathrm{s}\right)$, $n_{\text{dof }}^{\text{lower }}$ is controlled by the lower-body strategy $\pi_{\text{lower }}\left(a_{\text{lower }}\mid\mathrm{s}\right)$. The dual-agent reinforcement learning strategy is trained using the Proximal Policy Optimization (PPO) algorithm [32] and shares the same state space ${s}_{{t}}=\left[{s}_{{t}}^{\text{prop }},\mathrm{c}^{l},\mathrm{~g}^{{u}}\right]$, allowing both agents to consider information from the entire body and achieve coordinated control. Among them, the ${s}_{\mathrm{t}}^{\text{prop }}=\left[{q}_{{t}},\dot{q}_{t},\omega_{t},gv_{t},a_{t-1}\right]$ represents the body perception information of the robot, the $c^{l}=\left[\hat{v}_{x,t},\hat{v}_{y,t},\hat{\omega}_{yaw},\phi_{\text{stance}}\right]$ represents the command information for the lower-body strategy, and the ${g}^{{u}}\in\mathbb{R}^{14}$ represents the command information for the upper-body strategy. The details of the state and action Space are given in Appendix A.
The dual-agent policy is formulated as a dual-objective optimization problem. The value functions $\left(V^{l},V^{u}\right)$ corresponding to the upper and lower body policies $\left(\pi^{l},\pi^{u}\right)$, and the rewards obtained can be represented as $\left(r^{l}\left(s,a^{l},a^{u}\right),r^{u}\left(\mathrm{~s},\mathrm{a}^{1},\mathrm{a}^{\mathrm{u}}\right)\right)$. The specific reward functions and RL training parameters can be found in Appendix B and Appendix C. The dual policies are optimized by maximizing the cumulative rewards [29],
@@
Where $\gamma\in[0,1]$ represents the reward discount factor, $T$ denotes the total number of time steps, $\theta_{u}$ and $\theta_{l}$ respectively represent the policy parameters upper and lower bodies.
3.2 Robust lower-body locomotion under external forces
Robust locomotion of humanoid robots remains a significant challenge, primarily due to the complexity of the floating base and the instability caused by high inertia. The difficulty is further compounded when payloads are applied to the upper body, increasing the overall inertia and introducing additional modeling uncertainties, which complicate the control design.
For the lower-body strategy $\pi^{l}:C^{l}\times S\mapsto A^{l}$, various external disturbances are introduced during the training process to enhance the robot’s robustness, including adversarial upper-body perturbations and high-frequency random pushing disturbances. Additionally, we introduce various domain randomizations during training to improve the model’s generalization, with specific details provided in Appendix D.
3.3 Upper-body motion generation under external forces
The demonstration data for the upper-body policy are derived from the AMASS human-motion dataset [25] after feasibility filtering. Following PHC [24], human poses are retargeted to humanoid poses, and filtering is applied to suppress jitter and discontinuities. Instead of using open-loop replay, we train a reinforcement learning policy for the upper body that adapts to changing environments, avoids self-collisions, and improves whole-body coordination. The upper-body policy $\pi^{\mathrm{u}}:C^{{u}}\times S\mapsto A^{{u}}$ is trained in conjunction with the lower-body policy $\pi^{l}:C^{l}\times S\mapsto A^{l}$, and the state space of the upper-body strategy includes the state of the entire body. To enhance the load capacity of the humanoid robots, disturbances are applied to the robot’s end-effectors in the form of random 3D spatial forces [48]. Similar to the lower-body policy training, various domain randomizations are also incorporated during the upper-body policy training, with specific details provided in Appendix D.